# Autonomous Agents for Collaborative Task under Information Asymmetry

### v1

@@ Abstract
Large Language Model Multi-Agent Systems (LLM-MAS) have achieved great progress in solving complex tasks. It performs communication among agents within the system to collaboratively solve tasks, under the premise of shared information. However, when agents’ communication is leveraged to enhance human cooperation, a new challenge arises due to information asymmetry, since each agent can only access the information of its human user. Previous MAS struggle to complete tasks under this condition. To address this, we propose a new MAS paradigm termed iAgents, which denotes Informative Multi-Agent Systems. In iAgents, the human social network is mirrored in the agent network, where agents proactively exchange human information necessary for task resolution, thereby overcoming information asymmetry. iAgents employs a novel agent reasoning mechanism, InfoNav, to navigate agents’ communication towards effective information exchange. Together with InfoNav, iAgents organizes human information in a mixed memory to provide agents with accurate and comprehensive information for exchange. Additionally, we introduce InformativeBench, the first benchmark tailored for evaluating LLM agents’ task-solving ability under information asymmetry. Experimental results show that iAgents can collaborate within a social network of 140 individuals and 588 relationships, autonomously communicate over 30 turns, and retrieve information from nearly 70,000 messages to complete tasks within 3 minutes.
@@
—– Joey’s and Chandler’s agents’ discussion on the word "friend",
@@ 1 Introduction
There has been notable progress in autonomous agents driven by the Large Language Model (LLM) vaswani2017attention ; NEURIPS2020_1457c0d6 ; bubeck2023sparks ; touvron2023llama , especially in developing communicative agents for completing collaborative tasks xi2023rise ; wang2024survey ; guo2024large ; qian2023communicative ; hong2023metagpt ; tang2023medagents ; li2024agent ; zheng2023chatgpt , as shown in Figure 1a. In these multi-agent systems (MAS), multiple agents are created through role-play prompting li2024camel to imitate the ability of human experts and form a virtual entity (e.g., an agent company or hospital) to provide solutions derived from agents’ communication. Agents share context in the virtual entity to facilitate collective decision-making.
Given that autonomous communication among agents has achieved significant success in discussing, decomposing, and resolving various complex tasks, a natural idea is to introduce such autonomous communication into human society to enhance collaboration efficiency among humans. However, when we manage to assign each human with an agent, establishing a mapping from the human social network to the agent network so autonomous communication among agents can facilitate human cooperation, a new challenge arises. This challenge involves dealing with asymmetry Weber_1978 ; tomasello2009cultural in various types of information (environment, goals, and mind state) zhou2024sotopia ; zhou2024real ; premack1978does ; bubeck2023sparks ; xu2024opentom since each agent can only observe the information of its human user. Previous LLM-MAS are not suitable for handling this scenario, because 1) human information is sensitive and private, so the asymmetry can not be resolved by directly collecting all information into one place and sharing it as the context for MAS. 2) Human information is dynamic so it can not be easily memorized during pre-training and activated accurately through role-play prompting in MAS to avoid asymmetry. Essentially, agents’ cooperation in previous MAS has adopted an introspective approach within the virtual entity, which struggles to deal with asymmetry in human information.
To bridge gaps in such asymmetry, agents need to retrieve information from humans, and proactively exchange information with each other, creating a new ecosystem combining the human network and the agent network. Therefore, we propose the concept of iAgents (Informative Multi-Agent Systems) [^1] Available on https://github.com/thinkwee/iAgents for achieving this kind of collaboration, as shown in Figure 1b. iAgents utilizes a new method for agent reasoning (InfoNav) which models the agents’ minds and navigates agents’ autonomous communication toward proactive information exchange. Furthermore, a new memory mechanism is designed to provide agents with accurate and comprehensive information for exchange. Additionally, we introduced InformativeBench, the first benchmark evaluating agents’ collaboration in social networks with information asymmetry. It includes both information-seeking tasks for evaluating agents’ ability to resolve asymmetry in a large amount of information, and complex algorithm-like tasks that focus on evaluating agents’ ability to collaboratively reason under information asymmetry. We found that iAgents could perform effective communication and collaboration within a social network of (shown in Figure 5) 140 individuals and 588 relationships, and across over 30 dialogues they searched nearly 70,000 messages and resolved the task within 3 minutes. Despite such cases where iAgents exhibit impressive performance, agents with some state-of-the-art LLM backends achieved an average accuracy of 50.48% on InformativeBench, with the most challenging task achieving only 22.8% accuracy, which reveals both potential promise and challenges in this direction.
@@
Figure 1: Comparison between previous MAS and iAgents. The visibility range of information for each agent is highlighted with a colored background. In the previous MAS, all agents share all information (colored background of Virtual Company), so it only works under the premise of shared context. In iAgents, each agent only sees information of its human user (different colored backgrounds of cartoon characters), and iAgents is designed to deal with information asymmetry.
@@ 2 Method
Figure 2: Overall architecture of iAgents. From left to right, 1) each individual in the social network is equipped with an agent, and 2) two individuals summon their agents to solve a task, each initially holding the information that is visible to its human user. Then 3) agents automatically raise communication and exchange necessary information on behalf of human users. Finally, 4) agents perform a consensus check on their planning completed by InfoNav to solve the task.
@@ 2.2 Overview
To ensure each generated utterance provides valuable information and eliminates asymmetry, how to exchange information and what information to exchange is crucial. To deal with these two questions, we use the InfoNav mechanism to guide communication towards effective information exchange. Furthermore, we introduce the Mixed Memory mechanism which organizes human information into Fuzzy and Distinct Memory for accurate and comprehensive retrieval. Additionally, each agent can initiate new communication $C^{new}$ within their subnetwork, which means the communication $C$ may be recursive and can diffuse among the social network:
@@ 2.3 InfoNav
As shown in Equation 6, the agent needs to be aware of its rationale set and ongoing communication to effectively advance the conversation. While the status of the rationale set can be implicitly inferred from utterances, this inference is often unreliable for LLM Agents. This unreliability can lead to incorrect states and then generate meaningless utterances, such as repetitive questioning or redundant thanking, which makes it harder to infer rationale and creates a vicious cycle. To address this, we propose the InfoNav mechanism. InfoNav explicitly records the rationale set using the agent’s planning text and navigates the autonomous communication. Before each utterance, the agent reviews its plan to identify which unknown rationale to inquire about and then updates the plan based on the responses received. Figure 3 shows an example of InfoNav in action. Initially, we prompt the agent to generate a plan $P$ outlining the rationales needed to answer question $Q$. Since the agent has no information at the beginning, all rationales in the plan are marked as unknown:
@@
where $r^{u}$ denotes unknown rationales. During communication, these rationales can updated to "known" and filled with concrete content as information is provided. The plan is written in fluent natural language, making it explicit and effective for prompting the model. Therefore, using InfoNav, the rationale set $R$ in equation 6 is rewritten to plan $P$, and the updated rationale $R^{new}$ is replaced to the plan with filled rationales $P(r^{k})$, where $r^{k}$ represents known rationales:
@@
After multiple turns of communication, both sides finish the update of their plans. The agents then perform consensus reasoning, which unifies collected rationales and discards conflicting ones to reach an answer. Thus, equations 1 to 3 rewrite to:
@@ 2.4 Mixed Memory
In iAgents, agents are navigated by InfoNav to retrieve human information and share it with other agents for collaboration. Retrieval of human information is challenging due to it being diverse in format and complex to organize. We propose organizing human information into two types of agent memories: Distinct Memory and Fuzzy Memory. These memories facilitate reactive retrieval, ensuring accurate and comprehensive rationale extraction, as shown in Figure 2.
Distinct Memory ($Mem_{D}$) stores human information in a structured format, allowing for exact matches such as structured query language. Distinct memory faithfully preserves the original information ($I$) and supports accurate retrieval. Additionally, it enables information retrieval from multiple spans across different chat sessions ($s$), capturing evolving changes in rationales.
However, distinct memory’s strict exact-match requirements complicate the retrieval process. It also struggles to provide cohesive context. To address these issues, we introduce Fuzzy Memory ($Mem_{F}$). Fuzzy memory stores summarized session texts ($I_{s}$) and uses embedding-based ANN retrieval johnson2019billion . Although both fuzzy memory and reflection park2023generative produce more abstract text, we emphasize an objective summary of information to facilitate session-level retrieval, rather than subjective generalizations to aid in planning. While this approach may lose some details, it offers a comprehensive context and enables robust, semantic-based retrieval. Therefore, the retrieval action in Equation 8 can be rewritten to involve both memory types:
@@ 3.1 Pipelines
Currently, there exists no benchmark or dataset specifically designed to address information asymmetry in collaborative tasks among communicative agents. In this paper, we construct InformativeBench, the first benchmark to evaluate agent collaboration tasks featuring information asymmetry in social networks. It contains five datasets to assess agents’ various capabilities. What’s more, recent studies have found that LLM continuously ingests internet data so static benchmarks can be easily memorized and overfitted zhou2023don ; xu2024benchmarking ; zhang2024careful . Hence we share two pipelines for constructing InformativeBench which are easy to realize and can be generalized to more domains for constant and dynamic evaluations. They are Needle-Type and Algorithm-Type pipelines, as shown in Figure 4.
@@
Needle-Type Pipeline A "needle"kamradt2023needle is inserted into the social network, and agents are tasked with finding this "needle" information. This evaluates their ability to share and locate information. The dataset can be created either by splitting the needle and spreading it into the network, or by combining information pieces from the network to generate a needle. For the split method, the Needle in the Persona (NP) dataset modifies the dialogue in the SPC datasetjandaghi2023faithful by adding a common or opposite persona to two individuals’ personas. Agents are asked to find this persona. For the combination method, the FriendsTV dataset reconstructs the social network based on the entire season one script of Friends wiki:Friends_TV_series , involving 140 characters with 588 relationships (as shown in Figure 5), and combines two questions in the FriendsQA dataset yang2019friendsqa ; li2020transformers as "needle pieces" to generate new question. This dataset, the largest in InformativeBench, features sarcasm, plot twists, and complex relationships for simulating real-world challenges.
@@
Algorithm-Type Pipeline Humans are assigned different pieces of information, which serve as inputs for an algorithm. Agents must solve tasks where the answer is the algorithm’s output, evaluating their reasoning abilities based on information exchange. In InformativeBench, this is represented by the Schedule dataset, which develops a program for assigning different schedules to individuals. Agents are presented with algorithmic problems of varying difficulties, and the algorithm program automatically verifies the correctness of their solutions. The datasets include questions of three levels of difficulty: Easy) calculate the number of conflicting schedules between two people, Medium) find the longest activity among six people, and Hard) find the longest common free period among six people.
3.2 Question Distribution
Figure 6 presents the distribution of problem types in InformativeBench. It is evident that the majority of the questions in InformativeBench are of the "What" and "Who" types, which have objective ground truth and lack ambiguity. In the Schedule dataset, questions are categorized into three difficulty levels, with each difficulty level corresponding to a different type of question: "What", "How Many", and "How Long".
3.3 Question Sample
Dataset Question Sample
Needle in the Persona What fantasy series does Alice enjoy that Dave is indifferent about?
Table 1: Question sample in the InformativeBench.
Table 1 provides examples of problems from five datasets.
FriendsTV. Based on questions from the FriendsQA data, new questions are synthesized. For instance, as shown in Table 1, there is a scene in the third act of the seventh episode of the first season of Friends where a blackout occurs. Agents need to combine the rationales and answers to the original questions in FriendsQA, which are "Where did the blackout happen?" and "Who was worried about grandmother being affected by the blackout?", to locate this scene in the script of the first season and find the relevant characters.
3.4 Benchmark Statistic
Pipeline Needle Algorithm Algorithm Algorithm Needle
#QA 100 30 30 30 222
#Individuals 4 4 6 6 140
#Relationships 5 3 5 5 588
Need External Memory No No No No Yes
Metrics Precision Precision F1 IoU Precision
Table 2: Statistic of InformativeBench.

### v2

@@ Abstract
Large Language Model Multi-Agent Systems (LLM-MAS) have greatly progressed in solving complex tasks. It communicates among agents within the system to collaboratively solve tasks, under the premise of shared information. However, when agents’ collaborations are leveraged to perform multi-person tasks, a new challenge arises due to information asymmetry, since each agent can only access the information of its human user. Previous MAS struggle to complete tasks under this condition. To address this, we propose a new MAS paradigm termed iAgents, which denotes Informative Multi-Agent Systems. In iAgents, the human social network is mirrored in the agent network, where agents proactively exchange human information necessary for task resolution, thereby overcoming information asymmetry. iAgents employs a novel agent reasoning mechanism, InfoNav, to navigate agents’ communication towards effective information exchange. Together with InfoNav, iAgents organizes human information in a mixed memory to provide agents with accurate and comprehensive information for exchange. Additionally, we introduce InformativeBench, the first benchmark tailored for evaluating LLM agents’ task-solving ability under information asymmetry. Experimental results show that iAgents can collaborate within a social network of 140 individuals and 588 relationships, autonomously communicate over 30 turns, and retrieve information from nearly 70,000 messages to complete tasks within 3 minutes [^1] Available on https://github.com/thinkwee/iAgents..
@@
—– Joey’s and Chandler’s agents discuss the word "friend",
@@ 1 Introduction
Figure 1: Comparison between previous MAS (left) and iAgents (right). The visibility range of information for each agent is highlighted with a colored background. On the left, all agents share all information (colored background of Virtual Company). On the right, each agent could only see information about its human user (separated colored backgrounds), and iAgents is designed to deal with such kind of information asymmetry.
There has been notable progress in autonomous agents driven by the Large Language Model (LLM), especially in developing communicative agents for completing collaborative tasks xi2023rise ; wang2024survey ; guo2024large ; qian2023communicative ; hong2023metagpt ; tang2023medagents ; li2024agent ; zheng2023chatgpt ; qian-etal-2024-experiential , as shown in Figure 1a. In these multi-agent systems (MAS), multiple agents are created through role-play prompting li2024camel to imitate the ability of human experts and form a virtual entity (e.g., an agent company or hospital) to provide solutions derived from agents’ communication. Agents share context in the virtual entity to facilitate collective decision-making.
Since autonomous communication among agents has achieved significant success in discussing, decomposing, and resolving various complex tasks, the natural idea is to upgrade the tasks for agents from single-person to multi-person, where agents work on behalf of multiple human users and solve the collaboration task among these users. An intuitive solution is to assign each user an agent and perform autonomous collaborations among these agents. However, in such a setting, a new challenge arises. This challenge involves dealing with asymmetry Weber_1978 ; tomasello2009cultural in various types of information (environment, goals, and mind state) zhou2024sotopia ; zhou2024real ; premack1978does ; bubeck2023sparks ; xu2024opentom ; chen2024internet since each agent can only observe the information of its human user. Previous LLM-MAS are not suitable for handling this scenario, because 1) human information is sensitive and private, so the asymmetry can not be resolved by directly collecting all information into one place and sharing it as the context for MAS. 2) Human information is dynamic so it can not be easily memorized during pre-training and activated accurately through role-play prompting in MAS to avoid asymmetry. Essentially, agents’ cooperation in previous MAS has adopted an introspective approach within the virtual entity (an agent hospital/town/software company), which struggles to deal with asymmetry in human information.
To bridge gaps in such asymmetry, agents need to retrieve information from humans and proactively exchange information, creating a new ecosystem combining the human and the agent network. Therefore, we propose the concept of iAgents (Informative Multi-Agent Systems) for achieving this kind of collaboration, as shown in Figure 1b. iAgents utilizes a new agent reasoning method (InfoNav) to model the agents’ minds and navigate communication among agents toward proactive information exchange. Furthermore, a new memory mechanism is designed to provide agents with accurate and comprehensive information for exchange. Additionally, we introduced InformativeBench, the first benchmark evaluating agents’ collaboration ability under information asymmetry. It includes both information-seeking tasks within large social networks and algorithm-like reasoning tasks over a small network. Our contributions can be summarized as follows:
  1. 1.
    We raise the research problem of information asymmetry in the multi-agent system for enhancing human collaboration, which is the first to shift the research perspective in this area from a holistic system view to individuals within the system. It gives a new vision to the human-agent collaboration relationship.
  2. 2.
    We propose the iAgents framework to deal with the information asymmetry in a multi-agent system. Equipped with InfoNav and improved memory mechanism, iAgents could perform effective communication and collaboration within a social network (shown in Figure 6) of 140 individuals and 588 relationships, and across over 30 dialogues they searched nearly 70,000 messages and resolved the task within 3 minutes.
  3. 3.
    We introduce the first multi-agent information asymmetry benchmark, InformativeBench. Agents with some state-of-the-art LLM backends achieved an average accuracy of 50.48% on InformativeBench, with the most challenging task achieving only 22.8% accuracy, which reveals both potential promise and challenges in this direction.
2 Related Work
Agents based on Large Language Models (LLMs) Originate from ancient Greek philosophy, where an "agent" denoted a being capable of intentional action, driven by mental states such as desires and beliefs sep-agency . As AI progresses, this concept integrates into the simulation and understanding of intelligent behavior Wooldridge_Jennings_1995 . Traditionally, agent research focused on some specific tasks 10.1109/64.180407 ; 10.1145/176789.176804 ; sutton1999reinforcement , but the emergence of LLMs has shifted this focus. GPT-4, for instance, is recognized for achieving a form of general intelligence bubeck2023sparks , prompting exploration into equipping LLMs with agency and intrinsic motivation xi2023rise . Studies introduce frameworks for LLM-based agents, including memorisation zhang2024survey , decision-making, perception, and action modules, leveraging the inherent autonomy, reactivity, pro-activeness, and social ability of LLMs weng2023agent . Notable applications include utilizing single-agent systems and multi-agent systems for task solving and simulation qian2023communicative ; hong2023metagpt ; tang2023medagents ; li2024agent ; zheng2023chatgpt ; chen2024agentverse ; park2023generative ; qian2024iterative . However, most of this research has primarily focused on agent capabilities, often neglecting interaction and cooperation paradigms. This highlights a critical research gap boiko2023emergent , warranting further exploration in this area.
Human-Agent and Multi-Agent Cooperation Paradigms To ensure that agents align with human objectives kenton2021alignment ; paul2023 , human-agent cooperation is crucial. Two main paradigms of human-agent cooperation are the Equal-Partnership and the Instructor-Executor. The former emphasizes agents as communicators who understand human emotions Hasan_2023 , while the latter highlights the human’s guiding role, with agents following instructions gao2023assistgpt . However, single-agent systems face limitations, such as the inability to collaborate, learn from social interactions, and function effectively in complex scenarios Significant_Gravitas_AutoGPT ; liu2023training ; xi2023rise . Research suggests that multi-agent systems with each agent holding specific functions, can stimulate stronger intelligence minsky1988society ; Balaji2010AnIT ; qian2024scaling . Collaborative multi-agent systems can efficiently handle complex tasks mandi2023roco ; liang2023encouraging ; du2024multi . Recent studies focus on scenarios with information asymmetry among agents. For instance,  zhou2024sotopia explores social intelligence among agents achieving private goals based on common scene information. Additionally,  zhou2024real develops an evaluation framework to simulate social interactions with LLMs. The study finds that learning from omniscient simulations enhances interaction naturalness but doesn’t improve goal achievement in cooperative scenarios. Many real-world scenarios involve information asymmetry, posing challenges for multi-agent systems.
Reasoning In previous research, agents are tasked with providing accurate information to human users in human-machine collaboration. Due to the nature of language models, the output of information relies on the user’s input and the previously decoded content serving as rationale context. Therefore, some work on reasoning has explored how to improve the organization and expression of this rationale context NEURIPS2022_9d560961 ; NEURIPS2023_271db992 ; besta2024graph ; ding2023everything , to enhance the accuracy of output information. However, some research has also found that the reasoning process of LLMs differs from that of humans yang2024large ; jin2024impact ; chen2024masked ; pfau2024let . For questions relying on internal knowledge within LLMs to answer, agents do not necessarily solve problems step by step like humans. Instead, compared to steps, having context with sufficient information content is more important. For scenarios discussed in this paper, we focus on machine-to-machine communication, relying on external knowledge of LLMs to collaboratively answer questions. This poses different requirements for LLM reasoning, especially in terms of how to promote information flow so sufficient information is included in the context provided to LLM. Agents need to actively hu2024uncertainty and accurately acquire, provide, and ask for information.
@@ 3 Method
Figure 2: Overall architecture of iAgents. From left to right, 1) each individual in the social network is equipped with an agent, and 2) two human users invoke their agents to solve a task, each initially holding the information that is visible to its human user. Then 3) agents automatically raise communication and exchange necessary information on behalf of human users. Finally, 4) agents perform a consensus check on their planning completed by InfoNav to solve the task.
@@ 3.2 Overview
To ensure each generated utterance provides valuable information and eliminates asymmetry, how to exchange information and what information to exchange is crucial. To deal with these two questions, we use the InfoNav mechanism to guide communication towards effective information exchange. Furthermore, we introduce the Mixed Memory mechanism which organizes human information into Fuzzy and Clear Memory for accurate and comprehensive retrieval. Additionally, each agent can initiate new communication $C^{new}$ within their subnetwork, which means the communication $C$ may be recursive and can diffuse among the social network:
@@ 3.3 InfoNav
As shown in Equation 6, the agent needs to be aware of its rationale set and ongoing communication to effectively advance the conversation. While the status of the rationale set can be implicitly inferred from utterances, this inference is often unreliable for LLM Agents. This unreliability can lead to incorrect states and then generate meaningless utterances, such as repetitive questioning or redundant thanking, which makes it harder to infer rationale and creates a vicious cycle. To address this, we propose the InfoNav mechanism. InfoNav plans and tracks the status of the agent’s rationale set explicitly for better navigating the communication. Before each utterance, the agent reviews its plan to identify which unknown rationale to inquire about and then updates the plan based on the responses received. Figure 3 shows an example of InfoNav in action. Initially, we prompt the agent to generate a plan $P$ outlining the rationales needed to answer question $Q$. Since the agent has no information at the beginning, all rationales in the plan are marked as unknown:
@@
where $r^{u}$ denotes unknown rationales. During communication, if the agent gets the information of one rationale, it will update the status of this rationale from “unknown” to “known” and fill this information into the rationale placeholder in the planning text. The plan is written in fluent natural language, making it explicit and effective for prompting the model. Therefore, using InfoNav, the rationale set $R$ in equation 6 is rewritten to plan $P$, and the updated rationale $R^{new}$ is replaced to the plan with filled rationales $P(r^{k})$, where $r^{k}$ represents known rationales:
@@
After multiple turns of communication, both sides finish the update of their plans. Agents then unify collected rationales and discard conflicting ones to reach an answer, denoted as “Consensus Reasoning”. Thus, equations 1 to 3 rewrite to:
@@ 3.4 Mixed Memory
In iAgents, agents are navigated by InfoNav to retrieve human information and share it with other agents for collaboration. Retrieval of human information is necessary since 1) human’s lifelong information can not be stored in the “long context” (such as 128k tokens) of LLM, and 2) even though the information required for a single-turn conversation can fit into the context, the accumulation of information over multiple turns can lead to context explosion. It is also challenging to organize human information which is diverse in format and complex to understand. We propose organizing human information into two types of agent memories: Clear Memory and Fuzzy Memory. These memories facilitate reactive retrieval, ensuring accurate and comprehensive rationale extraction, as shown in Figure 2.
Clear Memory ($Mem_{C}$) stores information in a structured format to facilitate precise retrieval. Clear memory faithfully preserves the original information ($I$) and supports accurate retrieval. Additionally, it enables information retrieval from multiple spans across different chat sessions ($s$), capturing evolving changes in rationales.
However, Clear Memory’s strict exact-match requirements complicate the retrieval process. It also struggles to provide cohesive context. To address these issues, we introduce Fuzzy Memory ($Mem_{F}$). Fuzzy memory stores summarized session texts ($I_{s}$) and uses embedding-based ANN retrieval johnson2019billion . Although both fuzzy memory and reflection park2023generative produce summary-like text, we emphasize objective summarization of information to facilitate session-level retrieval, rather than subjective generalizations to aid in planning. While this approach may lose some details, it offers a comprehensive context and enables robust, semantic-based retrieval. Therefore, the retrieval action in Equation 8 can be rewritten to involve both memory types:
@@ 4 InformativeBench
To the best of our knowledge, there is no benchmark or dataset tailored for information asymmetry in the collaboration task among communicative agents. In this paper, we construct InformativeBench, the first benchmark to evaluate agent collaboration tasks featuring information asymmetry in social networks. It includes two categories with a total of five datasets. Details, including the scale, distribution, and metrics of the datasets, are provided in section C. What’s more, recent studies have found that LLM continuously ingests internet data so static benchmarks can be easily memorized and overfitted zhou2023don ; xu2024benchmarking ; zhang2024careful . Hence, two pipelines for constructing InformativeBench are easy to realize and can be generalized to more domains for constant and dynamic evaluations. They are Needle-Oriented and Reasoning-Oriented pipelines, as shown in Figure 4.
@@
Needle-Oriented Pipeline A "needle"kamradt2023needle is inserted into the social network, and agents are tasked with finding this "needle" information. This evaluates their ability to share and locate information. The dataset can be created by splitting the needle and spreading it into the network, or by collecting pieces from the network and combining them. For the split method, the Needle in the Persona (NP) dataset modifies the dialogue in the SPC datasetjandaghi2023faithful by adding a common or opposite persona to two individuals’ personas. Agents are asked to find this persona. For the combination method, the FriendsTV dataset reconstructs the social network from the entire Season 1 script of Friends wiki:Friends_TV_series , involving 140 characters with 588 relationships, and combines two questions in the FriendsQA dataset yang2019friendsqa ; li2020transformers as "needle pieces" to generate new question. This dataset, the largest in InformativeBench, features sarcasm, plot twists, and complex relationships for simulating real-world challenges.
@@
Reasoning-Oriented Pipeline Humans are assigned different pieces of information, which serve as inputs for an algorithm (such as sorting or merging). Agents must reason to get the answer which is the algorithm’s output. Therefore, the algorithm serves as an automatic verifier for information asymmetric reasoning. In InformativeBench, this is represented by the Schedule dataset, which develops a program for assigning different schedules to individuals. Agents are presented with algorithmic problems of varying difficulties, and the program automatically verifies the correctness of their solutions. The datasets include questions of three levels of difficulty: Easy) calculate the number of conflicting schedules between two people, Medium) find the longest activity among six people, and Hard) find the longest common free period among six people.